OpenStudy (anonymous):

Evaluate the definite integral. Use a graphing utility to verify your result. π/2 esin 3πx cos 3πx dx

OpenStudy (anonymous):

im trying to do u subsittion

OpenStudy (anonymous):

Are pi/2 and e the limits of integration?

OpenStudy (anonymous):

\[\int_?^{{\frac{\pi}2}}e^{\sin(3\pi x)}\cos(3\pi x)dx\]

OpenStudy (anonymous):

i am gong to bet that the lower limit is zero. that is just a guess. this would be an easy set up for u - sub

OpenStudy (anonymous):

no the limits of integration atre pi/2 and zero. the rest is the function to evalutate i did e^u and let u=sin3pix and then im stuck

OpenStudy (anonymous):

ah ok thought so

OpenStudy (anonymous):

you are in good shape.

OpenStudy (anonymous):

okay my s compute session is going ot end sooon OH NOW .

OpenStudy (anonymous):

okay my s compute session is going ot end sooon OH NOW .

OpenStudy (anonymous):

\[u=\sin(3\pi x)\] \[du=3\pi \cos(3\pi x)dx\]

OpenStudy (anonymous):

that aint is. try this one for a quick answer http://www.wolframalpha.com/input/?i=integral+e^%28sin%283*pi*x%29%29*cos%283*pi*x%29+dx+from+zero+to+pi%2F2

OpenStudy (anonymous):

change limits of integration \[u(0)=\sin(0)=0\] \[u(\frac{\pi}{2})=\sin(\frac{3\pi^2}{2})\] then \[\frac{1}{3\pi}\int_0^{\sin(\frac{3\pi^2}{2})}e^udu\]

OpenStudy (anonymous):

get \[\frac{e^{\sin(\frac{\pi^2}{2})}-1}{3\pi}\]

OpenStudy (anonymous):

thanks sooo much sorry i was at a public computer earlier and my session ended